A circle has a circumference of 100 meter. Find its area.
Answer by nerdybill(7384)

A circle has a circumference of 100 meter. Find its area.
.
circumference = 2(pi)r
100 = 2(pi)r
50 = (pi)r
50/(pi) = r
.
Area = (pi)r^2
Area = (pi)[50/(pi)]^2
Area = (pi)[50^2/(pi)^2]
Area = 50^2/(pi)
Area = 2500/(pi)
Area = 2500/(3.14)
Area = 796.18 square meters